CAUGHT BY WEB

■ learned about LP from a friend; ■ did adaptation of TUGBoat styles for GUST bulletin (had played with the experimental TEX mode of FWEB); ■ since then have been exploring the idea of LP;

THREE FACES OF WEB

■ LP programs are ‘human oriented’ – documentation is a part of LP program; ■ in code part you are telling to humans what computer should do, instead of telling to computer what should be done; ■ works well when writing programs directly from the keyboard;

UTILITARY BENEFITS OF WEB

■ LP programs have unique style; typeset LP programs are easy to read and comprehend; ■ code part could be read in a ‘stratified’ way which is quick and efficient way learning structure of programs; ■ indexes (automatically generated), cross-references, bibliography – everything typeset by TEX adds to readability

and understanding; ■ blends well with existing tools (RCS, AWK, PERL, DIFF, PATCH, MERGE...);

CREATING YOUR OWN LP TOOLS WITH NOWEB

■ TANGLE already done; ■ WEAVE modify existing scripts; ■ adapt existing format to work with macros inserted by the script; ■ extend tools;

WHAT I HAVE NOT SUCCEEDED TO TELL...

■ more about LP tools; ■ details of TEX-WEB system; ■ another approach for making changes / updates to LP sources;

TUG ’95 / 1